# Results of Fractional Rough Burgers equation in $H^{s}$ space and its application

### v1

@@ 2.3 Iteration argument
Assume that $f(x,y)$ satisfied some control property on $X$, such as
$\|f(x,y)\|_{X}\leq\|x\|_{X}+c\|y\|_{X},$ (2.6)
@@
Let $\mathcal{E}:=\dot{E}\times E$ be a Banach space, where $\dot{E}$ is subspace of $E$. Define norm space $\dot{E}\times E$ as
@@
Moreover define $\mathcal{B}$ a continuous bilinear map from $E\times E$ to $\dot{E}$, and $r$ a positive real number such that
$\|\mathcal{B}(f(x,y),f(x,y))\|_{\dot{E}}\leq\frac{1}{8r}(\|x_{n}\|_{\dot{E}}+c\|y_{n}\|_{E})^{2},$
where $f$ satisfied (2.5) and (2.6) for $X=E$. Let $\mathcal{G}(x)$ and $g(y)$ are linear map $E$ to $\dot{E}$ and satisfied
@@
Then there exist $\tilde{c}>0$, for any $(a,b)\in B(0,r)\subset\mathcal{E}$ and $0<c<\tilde{c}$, there exists unique $(y,x)\in B(0,4r)\subset\mathcal{E}$ such that
@@
We first prove that $(x_{n+1},y_{n+1})\in B(0,4r)\subset E$ when $(x_{n},y_{n})\in B(0,4r)\subset E$. In fact,
$\|(x_{n+1},y_{n+1})\|_{\dot{E}\times E}\leq$ $\|b\|_{E}+\|f(x_{n},y_{n})\|_{E}+2\|a\|_{\dot{E}}+2\|g(y_{n})\|_{\dot{E}}+2\|h(x_{n})\|_{\dot{E}}$
@@
$\leq$ $2\|a\|_{\dot{E}}+\|b\|_{E}+\frac{1}{8}\|y_{n}\|_{E}+\frac{1}{8}\|x_{n}\|_{\dot{E}}+\frac{1}{8}\|f(x_{n},y_{n})\|_{E}+\frac{1}{4r}(\|x_{n}\|_{\dot{E}}+c\|y_{n}\|_{E})^{2}$
$\leq$ $2\|a\|_{\dot{E}}+\|b\|_{E}+\frac{1}{8}\|y_{n}\|_{E}+\frac{1}{8}\|x_{n}\|_{\dot{E}}+\frac{1}{8}(\|x_{n}\|_{E}+c\|y_{n}\|_{E})+\frac{1}{4r}(\|x_{n}\|_{E}+c\|y_{n}\|_{E})^{2}$
$\leq$ $r+\frac{r}{2}+\frac{r}{2}+\frac{r}{4}+\frac{1}{8}(2r+4cr)+\frac{1}{8r}(4r^{2}+16cr^{2}+16c^{2}r^{2})<4r,$
for $0<c<\tilde{c}=\frac{1}{10}$. Denote $f_{n,m}=f(x_{n},y_{n})-f(x_{m},y_{m})\in E$, to prove $\{(x_{n},y_{n})\}_{n\in\mathbb{N}}$ is a Cauchy sequence, we estimate that
$\|(x_{n+1},y_{n+1})-(x_{n},y_{n})\|_{\dot{E}\times E}\leq$ $2\|x_{n+1}-x_{n}\|_{\dot{E}}+\|y_{n+1}-y_{n}\|_{E}$
$\leq$ $2\|\mathcal{G}(f_{n,n-1})\|_{\dot{E}}+2\|\mathcal{B}(f_{n,n-1},f(x_{n},y_{n}))+\mathcal{B}(f(x_{n},y_{n}),f_{n,n-1})\|_{\dot{E}}$
$+\|f_{n,n-1}\|_{E}$
$\leq$ $(\frac{7}{4}+c)\|x_{n}-x_{n-1}\|_{\dot{E}}+(c^{2}+\frac{13}{8}c+\frac{1}{8})\|y_{n}-y_{n-1}\|_{E}$
$<$ $\frac{37}{40}\|(x_{n},y_{n})-(x_{n-1},y_{n-1})\|_{\dot{E}\times E}.$
Then we get the limit of $(y_{n},x_{n})$ is a unique fixed point of mapping (2.7) in Banach space $\mathcal{E}$, which prove the lemma. ∎
@@ 3 Result in the range of $\gamma\in(\frac{4}{3},2]$
In this section, we consider the highly dissipative regime, which ensures that the linear component $X(t)$ possesses enhanced regularity properties. In fact, assume that $\gamma\in(\frac{4}{3},2]$, by Lemma 2.7, we have $X\in C_{T}H^{\frac{1}{6}+\delta}$ for some small $\delta$. It is subcritical since $X(t)\in H^{\frac{3}{2}-\gamma}.$
@@ 3.1 Bi-linear estimate
First we build a standard bilinearity estimate. Define $B(f,g)$ as following
$B(f,g)(t)=\int_{0}^{t}e^{(t-s)\Lambda^{\gamma}}\frac{1}{2}D(f(s)g(s))\mathrm{d}s,$ (3.1)
then we have following estimate.
@@
Fix $E_{T}:=L^{\infty}_{T}H^{s}$, with $0\leq\rho<\gamma-1$, $s=\sigma+\rho$, then we have
$\|B(f,g)\|_{E_{T}}\leq CT^{\frac{\gamma-1-\rho}{\gamma}}(1+T^{\frac{\rho}{\gamma}})\|fg\|_{H^{\sigma}}.$ (3.2)
@@
We write that
$\|B(f,g)(t)\|_{H^{s}}\lesssim$ $\int_{0}^{t}\|e^{-(t-\tau)\partial_{x}^{\gamma}}(1-\partial_{x}^{2})^{\frac{\rho}{2}}\partial_{x}(1-\partial_{x}^{2})^{\frac{\sigma}{2}}(fg)\|_{L^{2}}.$
We divide $\xi$ in two parts: $|\xi|\lesssim 1$ and $|\xi|\gtrsim 1$. For low frequencies $|\xi|\lesssim 1$, we have
$|e^{-(t-\tau)|\xi|^{\gamma}}(1+\xi^{2})^{\frac{\rho}{2}}\xi|\lesssim|e^{-(t-\tau)|\xi|^{\gamma}}|\xi||,$
For high frequencies $|\xi|\gtrsim 1$, we have
$|e^{-(t-\tau)|\xi|^{\gamma}}(1+\xi^{2})^{\frac{\rho}{2}}\xi|\lesssim|e^{-(t-\tau)|\xi|^{\gamma}}|\xi|^{\rho+1}|,$
Then we obtain that
$\|B(f,g)(t)\|_{H^{s}}\lesssim\int_{0}^{t}\max(\frac{1}{(t-\tau)^{\frac{1}{\gamma}}},\frac{1}{(t-\tau)^{\frac{1+\rho}{\gamma}}})\|fg\|_{H^{\sigma}}\lesssim T^{\frac{\gamma-1-\rho}{\gamma}}(1+T^{\frac{\rho}{\gamma}})\|fg\|_{H^{\sigma}},$
@@
The case of end-point $\rho=\gamma-1$ will be more complex. Using the time-frequencies analysis by frequencies localization, we need the estimate in $B_{2,\infty}^{s}$. In this paper, we don’t touch this endpoint. Since for any $\alpha<\frac{\gamma}{2}-\frac{1}{2}$, we can always have $X\in\mathcal{W}^{\alpha+}(T)$ a.s.
By Moser-type estimate, its not difficult to prove that
@@
Let $\gamma\in(\frac{4}{3},2]$, $\alpha\in(\frac{1}{6},\frac{1}{2})$, fix $2\delta=\alpha+\gamma-\frac{3}{2}$, then for $T\leq 1$ we have following bounds
@@
  • (2) For $f,g\in L^{\infty}_{T}W^{\alpha}$ for any $p\in[2,\infty]$, we have
@@
  • (3) For $f\in L_{T}^{\infty}{H^{\frac{1}{2}+\delta}}$ and $g\in L_{T}^{\infty}W^{\alpha}$ for any $p\in[2,\infty]$, then we have
@@ 3.2 Well-posedness
In this little section, we study the well-posedness of (1.1) for $\gamma>\frac{4}{3}$. Firstly, we consider the equation
$\partial_{t}v-\Lambda^{\gamma}v=\frac{1}{2}\partial_{x}(v^{2})+\partial_{x}(v\tilde{X})+\frac{1}{2}\partial_{x}(\tilde{X}^{2}),\quad u(0)=u_{0}.$ (3.3)
We solve the equation (3.3) by contraction mapping argument.
@@
Let $\alpha\in(\frac{1}{6},\frac{1}{2})$, $\gamma\in(\frac{4}{3},2]$ and $u_{0}\in H^{\frac{1}{2}+\delta}$ for $2\delta=\alpha+\gamma-\frac{3}{2}$. If $\tilde{X}(t)\in{C_{T}W^{\alpha}}$, then there exists a unique solution $v(t,x)$ of equation (3.3) with a postive $T^{*}$ satisfied $v\in C_{T^{*}}H^{\frac{1}{2}+\delta}$.
@@
Let $r=2\max(\|u_{0}\|_{H^{\alpha}},\|\tilde{X}\|_{C_{1}W^{\alpha}})$. We rewrite equation (3.3) as
$u(t,x)=P(t)u_{0}+B(v,\tilde{X})+B(\tilde{X},v)+B(\tilde{X},\tilde{X})+B(v,v).$
where $P(t)f$ satisfied linear equation
@@
By Lemma 2.8, we claim that there exist $T^{*}$ such that
$a:=P(T^{*})u_{0}+B(\tilde{X},\tilde{X})\in B(0,r),\quad\|B\|_{C_{T^{*}}H^{\frac{1}{2}+\delta}}\leq\frac{1}{8r},\quad\sup_{v}\frac{2\|B(v,\tilde{X})\|_{C_{T^{*}}H^{\frac{1}{2}+\delta}}}{\|v\|_{C_{T^{*}}H^{\frac{1}{2}+\delta}}}\leq\frac{1}{4}.$ (3.5)
For the first claim of (3.5), fix $T>0$ small enough to make sure that $C_{2}T^{\frac{\delta}{\gamma}}\|\tilde{X}\|^{2}_{C_{1}W^{\alpha}}\leq\frac{r}{4}$, then by Lemma 3.2, we have
@@
$\leq$ $\|u_{0}\|_{C_{T}H^{\frac{1}{2}+\delta}}+C_{2}T^{\frac{\delta}{\gamma}}\|\tilde{X}\|^{2}_{C_{1}W^{\alpha}}\leq r.$
For the second claim of (3.5), fix $T>0$ small enough to make sure that $C_{1}T^{\frac{\gamma-1}{\gamma}}\leq\frac{1}{8r}$, then by Lemma 3.2, we have
@@
$\leq$ $\frac{1}{8r}\|v\|_{C_{T}H^{\frac{1}{2}+\delta}}^{2}$
For the third claim of (3.5), fix $T$ small enough to make sure that $C_{3}T^{\frac{\delta}{\gamma}}\|\tilde{X}\|_{C_{1}W^{\alpha}}\leq\frac{1}{8}$, then by Lemma 3.2, we have
$\frac{\|B(v,\tilde{X})\|_{C_{T}H^{s}}}{\|v\|_{C_{T}H^{s}}}\leq$ $C_{3}T^{\frac{\delta}{\gamma}}\|\tilde{X}\|_{C_{1}W^{\alpha}}\leq\frac{1}{8}.$
Fix $T^{*}$ satisfied the condition above all, we finish our proof. ∎
@@
By Lemma 2.7, we have $X(t)\in C_{T}W^{\alpha}$ where $X(t)$ satisfied (1.5), by Proposition 3.1, we then we finish our proof by writing equation (1.1) as the (3.3) with fixing v=uXv=u-X and $\tilde{X}(t)=X(t)=\int_{0}^{t}P(t-s)\xi(s)\mathrm{d}s$. ∎
@@
We actually prove that the solution exists in the space $L_{t}^{\infty}H^{s}$ but depending on the equation itself, we can also easily demonstrate the continuity of $v$. Note that
$v(t)-v(s)=(S(t)-S(s))u_{0}+B_{s,t}(v+X,v+X)$
where $B_{s,t}(u,u):=\frac{1}{2}\int_{s}^{t}P(t-s)D(u^{2})\mathrm{d}s$. Use the similar argument for proving well-posedness, we can easily get the continuous.
To obtain the global result, we need more information of energy. Let us recall a classical energy estimate for general heat equation in [BCD11] in nonhomogeneous case and a energy control for Burgers equation.
@@
For $\gamma\in(\frac{3}{2},2]$, let $v$ be the smooth solution of (1.1) on $[0,T]$ for $u_{0}\in H^{\frac{1}{2}+\delta}$ and $X(t)$ satisfied (1.5). Then we have energy estimate as follows
$\|v\|_{L^{2}}^{2}+(1-\nu)\int_{0}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}\tau\lesssim\|u_{0}\|_{L^{2}}^{2}e^{C_{\nu}(1+\|X\|_{C_{t}W^{\alpha}}^{\frac{\gamma}{\gamma-1}})t}+C_{\nu}\int_{0}^{t}e^{C_{\nu}(1+\|X\|_{C_{t}W^{\alpha}}^{\frac{\gamma}{\gamma-1}})(t-t^{\prime})}\|X\|_{W^{\alpha}}^{4}\mathrm{d}t^{\prime}.$ (3.7)

### v2

@@ 2.3 Iteration argument
Assume that $f(x,y)$ satisfies some control property from $\dot{E}\times E$ to $X$, such as
$\|f(x,y)\|_{X}\leq\|x\|_{\dot{E}}+c\|y\|_{E},$ (2.6)
@@
Let $\dot{E},E,X$ be Banach spaces satisfying $\dot{E}\subset X\subset E$, $\mathcal{E}:=\dot{E}\times E$ be a Banach space with norm
@@
Moreover let $\mathcal{B}$ be a continuous bilinear map from $X\times X$ to $\dot{E}$, and $r$ a positive real number such that
$\|\mathcal{B}(u,v)\|_{\dot{E}}\leq\frac{1}{16r}\|u\|_{X}\|v\|_{X},$
Let $\mathcal{G}(x)$ and $g(y)$ be linear map from $E$ to $\dot{E}$ and satisfying
@@
Let $f$ satisfy (2.5) and (2.6) for $X$, then there exist $\tilde{c}>0$, such that for any $(a,b)\in B(0,\frac{r}{4})\subset\mathcal{E}$ and $0<c<\tilde{c}$, there exists unique $(x,y)\in B(0,4r)\subset\mathcal{E}$ such that
@@
We first prove that $(x_{n+1},y_{n+1})\in B(0,4r)\subset\mathcal{E}$ when $(x_{n},y_{n})\in B(0,4r)\subset\mathcal{E}$. Observing that $\|x_{n}\|\leq 2r$, $\|y_{n}\|\leq 4r$ and
$\|f(x_{n},y_{n})\|_{E}\leq\|f(x_{n},y_{n})\|_{X}\leq\|x_{n}\|_{\dot{E}}+c\|y_{n}\|_{E}\leq 2r+4cr,$
by a directly computation, we have
$\|(x_{n+1},y_{n+1})\|_{\mathcal{E}}\leq$ $\|b\|_{E}+\|f(x_{n},y_{n})\|_{E}+2\|a\|_{\dot{E}}+2\|g(y_{n})\|_{\dot{E}}+2\|h(x_{n})\|_{\dot{E}}$
@@
$\leq$ $2\|a\|_{\dot{E}}+\|b\|_{E}+\frac{1}{8}\|y_{n}\|_{E}+\frac{1}{8}\|x_{n}\|_{\dot{E}}+\frac{9}{8}\|f(x_{n},y_{n})\|_{E}+\frac{1}{8r}(\|f(x_{n},y_{n})\|_{X}^{2})$
$\leq$ $2\|a\|_{\dot{E}}+\|b\|_{E}+\frac{1}{8}\|y_{n}\|_{E}+\frac{1}{8}\|x_{n}\|_{\dot{E}}+\frac{9}{8}(\|x_{n}\|_{E}+c\|y_{n}\|_{E})+\frac{1}{8r}(\|x_{n}\|_{E}+c\|y_{n}\|_{E})^{2}$
$\leq$ $\frac{r}{4}+\frac{r}{2}+\frac{r}{4}+\frac{9}{8}(2r+4cr)+\frac{1}{8r}(4r^{2}+16cr^{2}+16c^{2}r^{2})$
$\leq$ $\frac{15}{4}r+(\frac{13}{2}c+2c^{2})r<4r,$
for $0<c<\tilde{c}=\frac{1}{30}$. Noting that $(x_{0},y_{0})\in B(0,\frac{r}{4})\subset B(0,4r)$, we obtain that $(x_{n},y_{n})\in B(0,4r)$ for all $n\in\mathbb{N}$. Setting $f_{n,m}=f(x_{n},y_{n})-f(x_{m},y_{m})\in E$, for all $n,m\in\mathbb{N}$, we have
$\|f_{n,m}\|_{X}\leq\|x_{n}-x_{m}\|_{\dot{E}}+c\|y_{n}-y_{m}\|_{E},\quad\|f(x_{n},y_{n})\|_{X}\leq\|x_{n}\|_{\dot{E}}+c\|y_{n}\|_{E}\leq 2r+4cr$
To prove that $\{(x_{n},y_{n})\}_{n\in\mathbb{N}}$ is a Cauchy sequence, we estimate that
$\|(x_{n+1},y_{n+1})-(x_{n},y_{n})\|_{\mathcal{E}}\leq$ $2\|x_{n+1}-x_{n}\|_{\dot{E}}+\|y_{n+1}-y_{n}\|_{E}$
$\leq$ $2\|g(y_{n}-y_{n-1})\|_{\dot{E}}+2\|h(x_{n}-x_{n-1})\|_{\dot{E}}+2\|\mathcal{G}(f_{n,n-1})\|_{\dot{E}}$
$+2\|\mathcal{B}(f_{n,n-1},f(x_{n},y_{n}))+\mathcal{B}(f(x_{n},y_{n}),f_{n,n-1})\|_{\dot{E}}+\|f_{n,n-1}\|_{E}$
$\leq$ $\frac{1}{8}\|y_{n}-y_{n-1}\|_{E}+\frac{1}{8}\|x_{n}-x_{n-1}\|_{\dot{E}}+\frac{1}{8r}(\|f_{n,n-1}\|_{X}\|f(x_{n},y_{n})\|_{X})$
$+\frac{1}{8r}(\|f_{n,n-1}\|_{X}\|f(x_{n-1},y_{n-1})\|_{X})+\frac{9}{8}\|f_{n,n-1}\|_{E}$
$\leq$ $(\frac{1}{8}+\frac{2cr+4c^{2}r}{4r}+\frac{9c}{8})\|y_{n}-y_{n-1}\|_{E}+(\frac{1}{8}+\frac{2r+4cr}{4r}+\frac{9}{8})\|x_{n}-x_{n-1}\|_{\dot{E}}$
$<$ $(\frac{7}{4}+c)\|(x_{n},y_{n})-(x_{n-1},y_{n-1})\|_{\mathcal{E}}<\frac{15}{16}\|(x_{n},y_{n})-(x_{n-1},y_{n-1})\|_{\mathcal{E}}$
Then we get the limit of $(x_{n},y_{n})$ is a unique fixed point of mapping (2.7) in Banach space $\mathcal{E}$, which prove the lemma. ∎
@@ 3 Result in the range of $\gamma\in(\frac{4}{3},2]$
This section treats the highly dissipative regime, in which the linear component $X(t)$ enjoys improved regularity. Specifically, for $\gamma\in(\frac{4}{3},2]$, Lemma 2.7 yields $X\in C_{T}H^{\frac{1}{6}+\delta}$ for a sufficiently small $\delta$. This regime is subcritical because the critical regularity for $X$ is $H^{\frac{3}{2}-\gamma}.$
@@ 3.1 Bi-linear estimate
First we build a standard bilinearity estimate. Defining $B(f,g)$ as follows
$B(f,g)(t)=\int_{0}^{t}e^{(t-s)\Lambda^{\gamma}}\partial_{x}(f(s)g(s))\mathrm{d}s,$ (3.1)
then we have the following estimate.
@@
Fixing $E_{T}:=L^{\infty}_{T}H^{s}$, with $0\leq\rho<\gamma-1$, $s=\sigma+\rho$, then we have
$\|B(f,g)\|_{E_{T}}\leq CT^{\frac{\gamma-1-\rho}{\gamma}}(1+T^{\frac{\rho}{\gamma}})\|fg\|_{C_{T}H^{\sigma}}.$ (3.2)
@@
By (3.1) and Parseval formula, we have
$\|B(f,g)(t)\|_{H^{s}}\lesssim$ $\int_{0}^{t}\|e^{-(t-\tau)\Lambda^{\gamma}}(1-\partial_{x}^{2})^{\frac{\rho}{2}}\partial_{x}(1-\partial_{x}^{2})^{\frac{\sigma}{2}}(fg)\|_{L^{2}}\mathrm{d}\tau$
$\lesssim$ $\int_{0}^{t}\|e^{-(t-\tau)|k|^{\gamma}}(1+|k|^{2})^{\frac{\rho}{2}}\partial_{x}(1+|k|^{2})^{\frac{\sigma}{2}}(fg)\|_{L^{2}}\mathrm{d}\tau$
Divide $k$ in two parts: $|k|\lesssim 1$ and $|k|\gtrsim 1$. For low frequencies $|k|\lesssim 1$, we have
$|e^{-(t-\tau)|k|^{\gamma}}(1+k^{2})^{\frac{\rho}{2}}k|\lesssim|e^{-(t-\tau)|k|^{\gamma}}|k||.$
For high frequencies $|k|\gtrsim 1$, we have
$|e^{-(t-\tau)|k|^{\gamma}}(1+k^{2})^{\frac{\rho}{2}}k|\lesssim|e^{-(t-\tau)|k|^{\gamma}}|k|^{\rho+1}|.$
Then for $t\in[0,T]$ we obtain that
$\|B(f,g)(t)\|_{H^{s}}\lesssim\int_{0}^{t}\max(\frac{1}{(t-\tau)^{\frac{1}{\gamma}}},\frac{1}{(t-\tau)^{\frac{1+\rho}{\gamma}}})\|fg\|_{H^{\sigma}}\mathrm{d}\tau\lesssim T^{\frac{\gamma-1-\rho}{\gamma}}(1+T^{\frac{\rho}{\gamma}})\|fg\|_{C_{T}H^{\sigma}},$
@@
The endpoint case $\rho=\gamma-1$ requires a more refined analysis, typically involving frequencies localization in Besov spaces $B_{2,\infty}^{s}$. Such endpoint issues are not addressed in the present work. Indeed, for any $\alpha<\frac{\gamma}{2}-\frac{1}{2}$, one has $X\in\mathcal{W}^{\alpha+}(T)$ almost surely, which suffices for our purposes.
By Moser-type estimate, it’s not difficult to prove that
@@
Let $\gamma\in(\frac{4}{3},2]$, $\alpha\in(\frac{1}{6},\frac{1}{2})$, $\delta$ satisfy $0<2\delta\leq\alpha+\gamma-\frac{3}{2}$, then for $T\leq 1$ we have the following bounds
@@
  • (2) For $f,g\in L^{\infty}_{T}W^{\alpha}$, we have
@@
  • (3) For $f\in L_{T}^{\infty}{H^{\frac{1}{2}+\delta}}$ and $g\in L_{T}^{\infty}W^{\alpha}$, then we have
@@ 3.2 Well-posedness
In this section, we study the well-posedness of equation (1.1) for $\gamma>\frac{4}{3}$. We begin by considering the equation
$\partial_{t}v-\Lambda^{\gamma}v=\partial_{x}(v^{2})+2\partial_{x}(v\tilde{X})+\partial_{x}(\tilde{X}^{2}),\quad u(0)=u_{0}.$ (3.3)
and solve it via a contraction mapping argument.
@@
Let $\alpha\in(\frac{1}{6},\frac{1}{2})$, $\gamma\in(\frac{4}{3},2]$ and $u_{0}\in H^{\frac{1}{2}+\delta}$ for $2\delta\leq\alpha+\gamma-\frac{3}{2}$. If $\tilde{X}(t)\in{C_{\bar{T}}W^{\alpha}}$ for some $\bar{T}>0$ , then there exists a unique solution $v(t,x)$ of equation (3.3) with a postive $T^{*}$ satisfying $v\in C_{T^{*}}H^{\frac{1}{2}+\delta}$.
@@
Let $r=2\max(\|u_{0}\|_{H^{\frac{1}{2}+\delta}},\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}})$. We rewrite equation (3.3) as
$v(t,x)=P(t)u_{0}+B(v,\tilde{X})+B(\tilde{X},v)+B(\tilde{X},\tilde{X})+B(v,v),$
where $P(t)f$ satisfying linear equation
@@
Firstly, we claim that there exist $0<T^{*}\leq 1\wedge\bar{T}$ such that for any $t\in[0,T^{*}]$
$a:=P(t)u_{0}+B(\tilde{X},\tilde{X})\in B(0,r),\quad\|B\|_{C_{t}H^{\frac{1}{2}+\delta}}\leq\frac{1}{8r},\quad\sup_{v}\frac{2\|B(v,\tilde{X})\|_{C_{t}H^{\frac{1}{2}+\delta}}}{\|v\|_{C_{t}H^{\frac{1}{2}+\delta}}}\leq\frac{1}{4}.$ (3.5)
For the first claim of (3.5), fixing $0<T\leq 1\wedge\bar{T}$ small enough to make sure that $C_{2}T^{\frac{\delta}{\gamma}}\|\tilde{X}\|^{2}_{C_{\bar{T}}W^{\alpha}}\leq\frac{r}{4}$, then by Lemma 3.2, we have
@@
$\leq$ $\|u_{0}\|_{C_{T}H^{\frac{1}{2}+\delta}}+C_{2}T^{\frac{\delta}{\gamma}}\|\tilde{X}\|^{2}_{C_{\bar{T}}W^{\alpha}}\leq r.$
For the second claim of (3.5), fixing $0<T\leq 1$ small enough to make sure that $C_{1}T^{\frac{\gamma-1}{\gamma}}\leq\frac{1}{8r}$, then by Lemma 3.2, we have
@@
$\leq$ $\frac{1}{8r}\|v\|_{C_{T}H^{\frac{1}{2}+\delta}}^{2}.$
For the third claim of (3.5), fixing $0<T\leq 1\wedge\bar{T}$ small enough to make sure that $C_{3}T^{\frac{\delta}{\gamma}}\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}\leq\frac{1}{8}$, then by Lemma 3.2, we have
$\frac{\|B(v,\tilde{X})\|_{C_{T}H^{s}}}{\|v\|_{C_{T}H^{s}}}\leq$ $C_{3}T^{\frac{\delta}{\gamma}}\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}\leq\frac{1}{8}.$
Let $T^{*}\leq\bar{T}$ satisfy all the condition above, we prove our claim is hold on $0,T^{*}]$. By Lemma 2.8, we finish our proof. ∎
@@
Let $\gamma\in(\frac{4}{3},2]$, by Lemma 2.7, we have $X(t)\in C_{T^{*}}W^{\alpha}$ for some $\alpha\in(\frac{1}{6},\frac{\gamma}{2}-\frac{1}{2})$ where $X(t)$ satisfies (1.5). By Proposition 3.1, then we finish our proof by writing equation (1.1) as the (3.3) with fixing v=uXv=u-X and $\tilde{X}(t)=X(t)=\int_{0}^{t}P(t-s)\xi(s)\mathrm{d}s$. ∎
@@
While we prove the existence of the solution in the space $L_{t}^{\infty}H^{s}$, the equation’s structure also allow the continuity of $v$ in time to be readily established. Observing that
$v(t)-v(s)=(S(t)-S(s))u_{0}+B_{s,t}(v+X,v+X),$
where $B_{s,t}(u,u):=\int_{s}^{t}P(t-s)D(u^{2})\mathrm{d}s$, Using the similar argument for proving well-posedness, we can easily get the continuous.
To obtain a global solution, further energy estimates are required. A classical energy estimate for the heat equation (see [BCD11]) is applied to derive the corresponding estimate for the rough Burgers equation.
@@
Let $\gamma\in(\frac{3}{2},2]$, $\alpha\in(\frac{1}{4},\frac{\gamma}{2}-\frac{1}{2})$, $\delta$ satisfy Theorem 1.1 and $0\leq T\leq\bar{T}<\infty$, let $v$ be the smooth solution of (3.3) on $[0,T]$ for $u_{0}\in H^{\frac{1}{2}+\delta}$ with $\tilde{X}\in C_{\bar{T}}W^{\alpha}$. Then for any $t\in[0,T]$, we have the following energy estimate
$\|v\|_{L^{2}}^{2}+(1-\nu)\int_{0}^{t}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq\|u_{0}\|_{L^{2}}^{2}e^{C_{\nu}(1+\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{\frac{2\gamma}{\gamma-1}})t}+C_{\nu}\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{4}\int_{0}^{t}e^{C_{\nu}(1+\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{\frac{2\gamma}{\gamma-1}})(t-t^{\prime})}\mathrm{d}t^{\prime}.$ (3.7)
for some small $\nu<1$.